# Knowledge Graphs for the Life Sciences: Recent Developments, Challenges and Opportunities

### v1

@@
Abstract.
The term life sciences refers to the disciplines that study living organisms and life processes, and include chemistry, biology, medicine, and a range of other related disciplines. Research efforts in life sciences are heavily data-driven, as they produce and consume vast amounts of scientific data, much of which is intrinsically relational and graph-structured.
The volume of data and the complexity of scientific concepts and relations referred to therein promote the application of advanced knowledge-driven technologies for managing and interpreting data, with the ultimate aim to advance scientific discovery.
In this survey and position paper, we discuss recent developments and advances in the use of graph-based technologies in life sciences and set out a vision for how these technologies will impact these fields into the future. We focus on three broad topics: the construction and management of Knowledge Graphs (KGs), the use of KGs and associated technologies in the discovery of new knowledge, and the use of KGs in artificial intelligence applications to support explanations (explainable AI). We select a few exemplary use cases for each topic, discuss the challenges and open research questions within these topics, and conclude with a perspective and outlook that summarizes the overarching challenges and their potential solutions as a guide for future research.
Key words and phrases:
@@ Abstract.
1. Introduction
The term life sciences refers to those disciplines that study living organisms and life processes, and include chemistry, biology, medicine, and a range of other related areas. Research efforts in life sciences are increasingly data-driven, as they produce and consume vast amounts of scientific data, much of which is intrinsically relational and graph-structured.
Much of this data is large-scale, complex, and presents many interrelationships and dependencies, thus being well suited to be represented in graph structures. For this reason, graph-based technologies are frequently used in the life sciences, and these disciplines have been drivers and early adopters of innovative methods and associated technologies.
In this brief survey and position paper we discuss recent developments and advances in the use of graph-based technologies in life sciences, and set out a vision for how these technologies will impact these fields in future. We illustrate the contribution in this paper in Figure 1.
We consider Knowledge Graphs (KGs) and their associated technologies to broadly include (i) different forms of graph-based representations, (ii) the logical languages that assign explicit semantics to such representations, and their associated automated reasoning technologies, and (iii) machine learning approaches that ingest data in graph-based representations and that process these graph-based representations to perform some task, e.g., data analytics.
These different forms of graph-based representations can be further categorized based on the type of content represented. We therefore distinguish schema-less and schema-based Knowledge Graphs. More specifically, a typical KG contains either or both a schema part (terminologies or TBox [^5] We introduce a list of key terms relevant to Knowledge Graphs and Life Sciences in Appendix A.) and a data part (facts, assertions, or ABox). The formal semantics of KGs can be expressed with the OWL ontology language [^6] Web Ontology Language: https://www.w3.org/OWL/.
In the remainder of this paper we will focus on three broad topic areas in which graph-based technologies have been used extensively, and we illustrate each area with some specific projects or use cases that guide our discussion and summary of the challenges that have been encountered.
@@ 1. Introduction
Figure 1. An overview illustration of definitions (upper right, in grey), topics (left column, in blue), use cases (middle), and challenges (bottom right, in green) for the research of KGs in the life sciences.
@@
We then provide a summary of the general challenges across the topics, that include intrinsic characteristics of KGs (e.g., scalability, evolution, heterogeneity) and their operational aspects in the real world (e.g., human interaction, personalization, distributed setting, and representation learning). We present the challenges by means of use cases and the current research efforts that address them. It is worth mentioning that while we aim to focus on the life sciences, many of the topics and challenges discussed in this work, especially those of KG construction and management in Section 3, are general and applicable to KGs in other domains such as finance, e-commerce, material, and urban management [114, 32], etc. The KG-based problem modeling and solving approaches in life science knowledge discovery could be applicable for addressing many other use cases and problems in a broader domain of AI for scientific discovery [178, 62].
In the next section, we introduce several different categories of KGs as they have been used in life sciences. Thereafter in Sections 3-5, each of the above topics is described in a dedicated section together with a survey of recent advances. Finally, in Section 6 we synthesize the overarching challenges and trends into a perspective on the outlook for the future.
2. Knowledge Graphs in the Life Sciences
KGs represent semantically-described real-world entities, typically through ontologies (vocabularies or schemas) [70, 63] and the data instantiating them, and thus provide descriptions of the entities of interest and their interrelations, by means of links to ontology classes describing them, organized in a graph [161]. KGs have been widely adopted in the life sciences, as can be seen in the composition of the Linked Open Data Cloud [^7] http://cas.lod-cloud.net, where life sciences represent one of the largest subdomains. A prominent example is the KG representing annotations regarding proteins by means of terms in the Gene Ontology describing different protein functions [4].
Whilst KGs are becoming increasingly popular in different domains including the life sciences, there is no single accepted definition of KG [45]. A KG can be formally described as a directed, edge-labeled graph $\mathcal{G}=(V,E)$, where $V$ refers to the vertices or nodes, representing real-world entities of interest (e.g., proteins, genes, compounds, cellular components, but also pathways, biological processes and molecular functions, to name a few) while $E$ refers to the edges in the graph, representing relationships or links between the entities in $V$ (e.g., binds, associates, etc.). These may be represented as statements about entities in the form of RDF [^8] Resource Description Framework: https://www.w3.org/RDF/ triples: (subject, predicate, object).
However, this formal definition only focuses on the components of KGs, but does not pose any constraint on what a KG should model or represent, and how. This is particularly true in life sciences, where the term Knowledge Graph has been used to refer to diverse graph data structures, typically interconnected, but often isolated.
@@ 2. Knowledge Graphs in the Life Sciences
Many of the everyday tasks faced by researchers in this domain require the systematic processing and integration of data and knowledge from data sources that are characterized by heterogeneous syntaxes and structures, formats, entity notation, schemas and scope, e.g., ranging from molecular mechanisms to phenotypes. Researchers in this area have been early adopters of Semantic Web and linked data approaches as a means to facilitate knowledge integration and processing to support tasks including semantic search, clinical decision support, enrichment analysis, data annotation and integration. However, a recent analysis of life science open data has identified several stand-alone data sources that exist in isolation, are not interlinked with other sources, and are schema-less (or use unpublished schemas), with limited reuse or mappings to other data sources [90]. Therefore, we can define a life sciences KG, following [133], as a data resource integrating one or more possibly curated sources of information into a graph whose nodes represent entities and edges represent relationships between two entities. This definition is consistent with other definitions found in the literature, e.g., [138].
These considerations underlie the reasons why KGs in life sciences can be of different types, and can be categorized across different dimensions. One of the most critical dimensions (in terms of support for complex queries and integration) is the categorization of KGs into schema-based and schema-less knowledge bases. In turn, the expressivity of the schema provides a further categorization criterion, depending on whether schemas are modelled as simple taxonomies (e.g., the NCBI taxonomy [157] included in the UMLS Metathesaurus [10]), RDFS [^9] RDF Schema: https://www.w3.org/TR/rdf-schema/ vocabularies or (fully axiomatized) OWL ontologies. In particular, this paper refers to this broad definition of KGs, which we then divide into:
@@
3. Knowledge Graph Construction and Management
The adoption of KGs in the life sciences is motivated by the need for standardisation of taxonomies and vocabularies to support the integration, exchange and analysis of data. More recently, richly annotated data is also being used in combination with machine learning methods for many applications, including helping to overcome issues related to the sparsity of data and helping to select promising candidates for reducing expensive and time-consuming physical experiments [66]. Graph-based machine learning approaches such as Graph Neural Networks have been applied to a number of life science tasks [51], including drug repurposing [123] and predicting polypharmacy side effects [199].
Given the diverse nature of the knowledge and tasks supported by KGs, the focus of state-of-the-art approaches has been the description of how individual KGs are developed within the specific domain [193], typically in terms of the specific approaches used for the development of the KG (e.g., data extraction process, relation extraction and entity discovery), rather than on the overall development process. More recently, some efforts have focused on providing an overview of development approaches and pipelines for the construction of KGs in the life sciences, and beyond [133, 167]. The process of constructing a KG depends heavily on:
@@ 3. Knowledge Graph Construction and Management
3.1. Knowledge Graph Construction Phases
@@
3.1.1. Data source selection
@@
3.1.2. Ontology construction
The aim of this phase is to define a common, consensus-based, controlled vocabulary to describe the data in an ontology [149]. The existence of a common structure, or schema, supports querying, integration and reasoning tasks over the KG.
@@ 3.1.2. Ontology construction
Traditional ontology engineering approaches are divided into top-down or bottom-up. Top-down approaches are based on more or less formal ontology engineering methodologies [47, 98, 134] or common practices [3] to build ontologies from a description of the domain elicited from domain experts [132], and/or by reusing or extending existing ontologies [84]. Ontology engineering methodologies define the ontology development process in terms of requirement analysis, entity and property definitions, ontology reuse, validation and population. In contrast, bottom-up approaches utilize semi-automatic data driven techniques, e.g., ontology learning from text [113], and can be used to refine and validate an ontology. These approaches are discussed in more detail when presenting the knowledge ingestion and validation phase.
Whilst general purpose ontology engineering methodologies have evolved to be used in the development of KGs [142], a considerable number of ontologies in the life science domain have been built as part of the Open Biological and Biomedical Ontologies (OBO) Foundry effort, [^14] https://obofoundry.org which defines a set of development principles for biological and biomedical ontologies and provides a suite of high-quality, interoperable, free and open source tools that support ontology development [118].
3.1.3. Knowledge extraction
@@
Knowledge extraction refers to the identification of entities and their relations from the data sources, which is a crucial step in the development of a KG [167]. Entity extraction identifies entities from the various data sources selected using Natural Language Processing (NLP) approaches and text mining techniques to analyse and extract relevant information from large text corpora [181, 106, 73]. Named entity recognition (NER) supports the identification of named entities in text, such as drug names, diseases, or chemical compounds, and their classification according to pre-defined entity types [130]. NER approaches in the life sciences are typically based on labour intensive tasks such as the definition of generic (e.g., orthographic, morphological, or dictionary-based) and specific rules that are typically defined by experts, and are not easily applicable to other corpora [198]. There are a number of issues hindering these approaches: a) the pace of scientific discovery and the identification of new entities; b) the large number of synonyms and term variations associated with an entity; and c) entity identifiers that are composed of a mixture of letters, symbols and punctuation, often in large sentences [104]. More recent approaches have proposed the use of supervised machine learning methods (e.g., conditional random fields, or Support Vector Machines, SVMs, neural networks, and neural language models in particular) [115, 88, 36] either in isolation, or combined in hybrid approaches to improve accuracy [152].
Entity recognition generates entities that are isolated and not linked [167]. The goal of Relation extraction is to discover relationships of interest between a pair of entities, thus describing their interaction. Relation extraction is a necessary step for entities defined in semi-structured or unstructured sources, whereas structured data sources are characterized by explicitly identifiable relationships. Typical approaches for relation extraction include rule-based [77, 148, 147], supervised [109, 50] and unsupervised approaches [101, 133]. Rule-based relation extraction identifies keywords (based on existing ontologies or expert defined dictionaries) and grammatical patterns to discover relations between entities. Supervised relationship extraction methods utilize publicly available pre-labelled datasets (e.g., BioInfer [144] or BioCreative II [100]) to construct generalized patterns that separate positive examples (sentences implying the existence of a relationship) from negative ones. Supervised approaches include SVMs, Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs) [7, 133]. Unsupervised relation extraction methods [116] have emerged to address the lack of scalability of supervised relation extraction methods, due to the high cost of human annotation. Unsupervised methods involve some form of clustering or statistical computation to detect the co-occurrence of two entities in the same text [133].
More recently, end-to-end approaches (End-to-End Relation Extraction – RE) have been used to tackle both tasks simultaneously. In this scenario, a model is trained simultaneously on both the NER and Relation Extraction objectives [76]. Furthermore, rule-based approaches can be combined with relation classification using specialized pre-trained language models adapted for life science domains, e.g., BioBERT [105], SapBERT [111], and RoBERTa-PM [107], to name a few. There is also a recent trend to probe and prompt pre-trained language models to extract relations (e.g., disease-to-disease, disease-to-symptoms) [190, 166].
3.1.4. Knowledge ingestion and validation
The aim of this phase is to ingest the entities and relationships extracted in a previous phase, which models knowledge from different sources. These entities and relations can be incomplete, ambiguous or redundant, and need to be appropriately aligned and integrated, and finally annotated according to the ontology constructed in phase 2.
Knowledge integration or fusion can critically improve the quality of data by performing entity resolution, i.e., the detection of different descriptions of the same real-world entity (also called entity matching, deduplication, entity linkage or entity canonicalization), prior to ingesting them in the KG. This reconciliation step is particularly crucial in the life sciences, where duplication can be caused by data modelled using different vocabularies or ontologies, or when data is extracted from literature sources that are rapidly changing. The severity of the ambiguity depends on the number of ontologies available for the domain. For instance, the number of gene vocabularies is far smaller than the number of disease vocabularies that could be present in the ingested datasets. Linking these entities requires costly alignment processing; in particular the alignment of disease entities is especially problematic given the number of different coding systems, whose conversion is often not trivial [53]. We further explore this issue in two of the use cases presented in Section 3.3, where we explore the problem of aligning vocabularies and ontologies through the use of mapping repositories and instance matching in automated clinical coding.
Entities are assigned unique identifiers (URI or IRI) that support the definition of bespoke namespaces, and support integration by reusing identifiers in related namespaces. Entity resolution is based on clustering similar entities together in a block, where similarity measures are used to detect duplicates [167]. Typical methods include sorted neighborhoods and traditional blocking; and machine learning methods are commonly used for similarity computation, e.g., feature vector computation [96].
This phase may also include the bottom-up construction of the ontology for those applications where a top-down approach is not feasible. Bottom-up approaches extract the relevant knowledge first, and then they construct the data schema / ontology based on the extracted data, typically using (semi-)automated methods, based on machine learning. Ontologies define the structure of the knowledge graph, which supports querying and data analytics. In bottom-up ontology development the structure of the knowledge graph is determined based on the extracted knowledge, thus providing a structure for this knowledge [71].
Often the construction of ontologies (either bottom-up or top-down) relies on the ability to correctly align and reuse entities defined across different domains and KGs. Furthermore, reuse of (or conformance to) existing upper level ontologies, e.g., BFO (Basic Formal Ontology) [3] provides the basis for the consistent and unambiguous formal definition of entities and relations that prevents errors in coding and annotation. The alignment of ontologies in life sciences and other domains is an active area of research, and we provide an overview of recent technical developments and challenges in Section 3.3.
Whilst bottom-up approaches, especially those based on alignment, are becoming more viable, especially given the support of language models, such as BERT [65], their performance is not always adequate for the task, as discussed in the second challenge in Section 3.4.
Knowledge enrichment and completion improve the KG quality by performing reasoning (KG materialization), inference [58] and optimization. Reasoning and inference support the assertion of new relations based either on logical reasoning (e.g., [131, 173]) or machine learning techniques (e.g., statistical relational learning or through embedding based link predictors for new concepts [35, 36, 68, 78] and node classifiers, also called KG refinement [138]). The extent and type of logical inferences depends on the expressivity of the ontology built in phase 2, or in a bottom-up fashion in this phase, together with any associated mappings. Description Logic formalisms, such as OWL, use logic-based reasoning for detecting and correcting incorrect assertions and ontology alignments [25].
3.1.5. KG storage and inspection

### v4

@@
Abstract
The term life sciences refers to the disciplines that study living organisms and life processes, and include chemistry, biology, medicine, and a range of other related disciplines. Research efforts in life sciences are heavily data-driven, as they produce and consume vast amounts of scientific data, much of which is intrinsically relational and graph-structured. The volume of data and the complexity of scientific concepts and relations referred to therein promote the application of advanced knowledge-driven technologies for managing and interpreting data, with the ultimate aim to advance scientific discovery. In this survey and position paper, we discuss recent developments and advances in the use of graph-based technologies in life sciences and set out a vision for how these technologies will impact these fields into the future. We focus on three broad topics: the construction and management of Knowledge Graphs (KGs), the use of KGs and associated technologies in the discovery of new knowledge, and the use of KGs in artificial intelligence applications to support explanations (explainable AI). We select a few exemplary use cases for each topic, discuss the challenges and open research questions within these topics, and conclude with a perspective and outlook that summarizes the overarching challenges and their potential solutions as a guide for future research.
keywords:
@@ Abstract
1 Introduction
The term life sciences refers to those disciplines that study living organisms and life processes, and include chemistry, biology, medicine, and a range of other related areas. Research efforts in life sciences are increasingly data-driven, as they produce and consume vast amounts of scientific data, much of which is intrinsically relational and graph-structured. Much of this data is large-scale, complex, and presents many interrelationships and dependencies, thus being well suited to be represented in graph structures. For this reason, graph-based technologies are frequently used in the life sciences, and these disciplines have been drivers and early adopters of innovative methods and associated technologies. In this brief survey and position paper we discuss recent developments and advances in the use of graph-based technologies in life sciences, and set out a vision for how these technologies will impact these fields in future. We illustrate the contribution in this paper in Figure 1. We consider Knowledge Graphs (KGs) and their associated technologies to broadly include (i) different forms of graph-based representations, (ii) the logical languages that assign explicit semantics to such representations, and their associated automated reasoning technologies, and (iii) machine learning approaches that ingest data in graph-based representations and that process these graph-based representations to perform some task, e.g., data analytics. These different forms of graph-based representations can be further categorized based on the type of content represented. We therefore distinguish schema-less and schema-based Knowledge Graphs. More specifically, a typical KG contains either or both a schema part (terminologies or TBox [^2] We introduce a list of key terms relevant to Knowledge Graphs and Life Sciences in Appendix A.) and a data part (facts, assertions, or ABox). The formal semantics of KGs can be expressed with the OWL ontology language [^3] Web Ontology Language: https://www.w3.org/OWL/. In the remainder of this paper we will focus on three broad topic areas in which graph-based technologies have been used extensively, and we illustrate each area with some specific projects or use cases that guide our discussion and summary of the challenges that have been encountered.
@@ 1 Introduction
Figure 1: An overview illustration of definitions (upper right, in gray), topics (left column, in blue), use cases (middle), and challenges (bottom right, in green) for the research of KGs in the life sciences.
@@
We then provide a summary of the general challenges across the topics, that include intrinsic characteristics of KGs (e.g., scalability, evolution, heterogeneity) and their operational aspects in the real world (e.g., human interaction, personalization, distributed setting, and representation learning). We present the challenges by means of use cases and the current research efforts that address them. It is worth mentioning that while we aim to focus on the life sciences, many of the topics and challenges discussed in this work, especially those of KG construction and management in Section 3, are general and applicable to KGs in other domains such as finance, e-commerce, material, and urban management [111, 31], etc. The KG-based problem modeling and solving approaches in life science knowledge discovery could be applicable for addressing many other use cases and problems in a broader domain of AI for scientific discovery [175, 60]. In the next section, we introduce several different categories of KGs as they have been used in life sciences. Thereafter in Sections 3-5, each of the above topics is described in a dedicated section together with a survey of recent advances. Finally, in Section 6 we synthesize the overarching challenges and trends into a perspective on the outlook for the future.
2 Knowledge Graphs in the Life Sciences
@@
KGs represent semantically-described real-world entities, typically through ontologies (vocabularies or schemas) [68, 61] and the data instantiating them, and thus provide descriptions of the entities of interest and their interrelations, by means of links to ontology classes describing them, organized in a graph [158]. KGs have been widely adopted in the life sciences, as can be seen in the composition of the Linked Open Data Cloud [^4] http://cas.lod-cloud.net, where life sciences represent one of the largest subdomains. A prominent example is the KG representing annotations regarding proteins by means of terms in the Gene Ontology describing different protein functions [4]. Whilst KGs are becoming increasingly popular in different domains including the life sciences, there is no single accepted definition of KG [43]. A KG can be formally described as a directed, edge-labeled graph $\mathcal{G}=(V,E)$, where $V$ refers to the vertices or nodes, representing real-world entities of interest (e.g., proteins, genes, compounds, cellular components, but also pathways, biological processes and molecular functions, to name a few) while $E$ refers to the edges in the graph, representing relationships or links between the entities in $V$ (e.g., binds, associates, etc.). These may be represented as statements about entities in the form of RDF [^5] Resource Description Framework: https://www.w3.org/RDF/ triples: (subject, predicate, object). However, this formal definition only focuses on the components of KGs, but does not pose any constraint on what a KG should model or represent, and how. This is particularly true in life sciences, where the term Knowledge Graph has been used to refer to diverse graph data structures, typically interconnected, but often isolated. Many of the everyday tasks faced by researchers in this domain require the systematic processing and integration of data and knowledge from data sources that are characterized by heterogeneous syntaxes and structures, formats, entity notation, schemas and scope, e.g., ranging from molecular mechanisms to phenotypes. Researchers in this area have been early adopters of Semantic Web and linked data approaches as a means to facilitate knowledge integration and processing to support tasks including semantic search, clinical decision support, enrichment analysis, data annotation and integration. However, a recent analysis of life science open data has identified several stand-alone data sources that exist in isolation, are not interlinked with other sources, and are schema-less (or use unpublished schemas), with limited reuse or mappings to other data sources [88]. Therefore, we can define a life sciences KG, following [130], as a data resource integrating one or more possibly curated sources of information into a graph whose nodes represent entities and edges represent relationships between two entities. This definition is consistent with other definitions found in the literature, e.g., [135]. These considerations underlie the reasons why KGs in life sciences can be of different types, and can be categorized across different dimensions. One of the most critical dimensions (in terms of support for complex queries and integration) is the categorization of KGs into schema-based and schema-less knowledge bases. In turn, the expressivity of the schema provides a further categorization criterion, depending on whether schemas are modeled as simple taxonomies (e.g., the NCBI taxonomy [154] included in the UMLS Metathesaurus [9]), RDFS [^6] RDF Schema: https://www.w3.org/TR/rdf-schema/ vocabularies or (fully axiomatized) OWL ontologies. In particular, this paper refers to this broad definition of KGs, which we then divide into:
@@
3 Knowledge Graph Construction and Management
The adoption of KGs in the life sciences is motivated by the need for standardization of taxonomies and vocabularies to support the integration, exchange and analysis of data. More recently, richly annotated data is also being used in combination with machine learning methods for many applications, including helping to overcome issues related to the sparsity of data and helping to select promising candidates for reducing expensive and time-consuming physical experiments [64]. Graph-based machine learning approaches such as Graph Neural Networks have been applied to a number of life science tasks [49], including drug repurposing [120] and predicting polypharmacy side effects [196]. Given the diverse nature of the knowledge and tasks supported by KGs, the focus of state-of-the-art approaches has been the description of how individual KGs are developed within the specific domain [190], typically in terms of the specific approaches used for the development of the KG (e.g., data extraction process, relation extraction and entity discovery), rather than on the overall development process. More recently, some efforts have focused on providing an overview of development approaches and pipelines for the construction of KGs in the life sciences, and beyond [130, 164]. The process of constructing a KG depends heavily on:
@@ 3 Knowledge Graph Construction and Management
3.1 Knowledge Graph Construction Phases
@@
3.1.1 Data source selection
@@
3.1.2 Ontology construction
@@
The aim of this phase is to define a common, consensus-based, controlled vocabulary to describe the data in an ontology [146]. The existence of a common structure, or schema, supports querying, integration and reasoning tasks over the KG. Traditional ontology engineering approaches are divided into top-down or bottom-up. Top-down approaches are based on more or less formal ontology engineering methodologies [45, 95, 131] or common practices [3] to build ontologies from a description of the domain elicited from domain experts [129], and/or by reusing or extending existing ontologies [82]. Ontology engineering methodologies define the ontology development process in terms of requirement analysis, entity and property definitions, ontology reuse, validation and population. In contrast, bottom-up approaches utilize semi-automatic data driven techniques, e.g., ontology learning from text [110], and can be used to refine and validate an ontology. These approaches are discussed in more detail when presenting the knowledge ingestion and validation phase. Whilst general purpose ontology engineering methodologies have evolved to be used in the development of KGs [139], a considerable number of ontologies in the life science domain have been built as part of the Open Biological and Biomedical Ontologies (OBO) Foundry effort, [^11] https://obofoundry.org which defines a set of development principles for biological and biomedical ontologies and provides a suite of high-quality, interoperable, free and open source tools that support ontology development [115].
3.1.3 Knowledge extraction
@@
Knowledge extraction refers to the identification of entities and their relations from the data sources, which is a crucial step in the development of a KG [164]. Entity extraction identifies entities from the various data sources selected using Natural Language Processing (NLP) approaches and text mining techniques to analyze and extract relevant information from large text corpora [178, 103, 71]. Named entity recognition (NER) supports the identification of named entities in text, such as drug names, diseases, or chemical compounds, and their classification according to pre-defined entity types [127]. NER approaches in the life sciences are typically based on labor intensive tasks such as the definition of generic (e.g., orthographic, morphological, or dictionary-based) and specific rules that are typically defined by experts, and are not easily applicable to other corpora [195]. There are a number of issues hindering these approaches: a) the pace of scientific discovery and the identification of new entities; b) the large number of synonyms and term variations associated with an entity; and c) entity identifiers that are composed of a mixture of letters, symbols and punctuation, often in large sentences [101]. More recent approaches have proposed the use of supervised machine learning methods (e.g., conditional random fields, or Support Vector Machines, SVMs, neural networks, and neural language models in particular) [112, 86, 35] either in isolation, or combined in hybrid approaches to improve accuracy [149]. Entity recognition generates entities that are isolated and not linked [164]. The goal of Relation extraction is to discover relationships of interest between a pair of entities, thus describing their interaction. Relation extraction is a necessary step for entities defined in semi-structured or unstructured sources, whereas structured data sources are characterized by explicitly identifiable relationships. Typical approaches for relation extraction include rule-based [75, 145, 144], supervised [106, 48] and unsupervised approaches [98, 130]. Rule-based relation extraction identifies keywords (based on existing ontologies or expert defined dictionaries) and grammatical patterns to discover relations between entities. Supervised relationship extraction methods utilize publicly available pre-labelled datasets (e.g., BioInfer [141] or BioCreative II [97]) to construct generalized patterns that separate positive examples (sentences implying the existence of a relationship) from negative ones. Supervised approaches include SVMs, Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs) [6, 130]. Unsupervised relation extraction methods [113] have emerged to address the lack of scalability of supervised relation extraction methods, due to the high cost of human annotation. Unsupervised methods involve some form of clustering or statistical computation to detect the co-occurrence of two entities in the same text [130]. More recently, end-to-end approaches (End-to-End Relation Extraction – RE) have been used to tackle both tasks simultaneously. In this scenario, a model is trained simultaneously on both the NER and Relation Extraction objectives [74]. Furthermore, rule-based approaches can be combined with relation classification using specialized pre-trained language models adapted for life science domains, e.g., BioBERT [102], SapBERT [108], and RoBERTa-PM [104], to name a few. There is also a recent trend to probe and prompt pre-trained language models to extract relations (e.g., disease-to-disease, disease-to-symptoms) [187, 163].
3.1.4 Knowledge ingestion and validation
The aim of this phase is to ingest the entities and relationships extracted in a previous phase, which models knowledge from different sources. These entities and relations can be incomplete, ambiguous or redundant, and need to be appropriately aligned and integrated, and finally annotated according to the ontology constructed in phase 2. Knowledge integration or fusion can critically improve the quality of data by performing entity resolution, i.e., the detection of different descriptions of the same real-world entity (also called entity matching, deduplication, entity linkage or entity canonicalization), prior to ingesting them in the KG. This reconciliation step is particularly crucial in the life sciences, where duplication can be caused by data modeled using different vocabularies or ontologies, or when data is extracted from literature sources that are rapidly changing. The severity of the ambiguity depends on the number of ontologies available for the domain. For instance, the number of gene vocabularies is far smaller than the number of disease vocabularies that could be present in the ingested datasets. Linking these entities requires costly alignment processing; in particular the alignment of disease entities is especially problematic given the number of different coding systems, whose conversion is often not trivial [51]. We further explore this issue in two of the use cases presented in Section 3.3, where we explore the problem of aligning vocabularies and ontologies through the use of mapping repositories and instance matching in automated clinical coding. Entities are assigned unique identifiers (URI or IRI) that support the definition of bespoke namespaces, and support integration by reusing identifiers in related namespaces. Entity resolution is based on clustering similar entities together in a block, where similarity measures are used to detect duplicates [164]. Typical methods include sorted neighborhoods and traditional blocking; and machine learning methods are commonly used for similarity computation, e.g., feature vector computation [93]. This phase may also include the bottom-up construction of the ontology for those applications where a top-down approach is not feasible. Bottom-up approaches extract the relevant knowledge first, and then they construct the data schema / ontology based on the extracted data, typically using (semi-)automated methods, based on machine learning. Ontologies define the structure of the knowledge graph, which supports querying and data analytics. In bottom-up ontology development the structure of the knowledge graph is determined based on the extracted knowledge, thus providing a structure for this knowledge [69]. Often the construction of ontologies (either bottom-up or top-down) relies on the ability to correctly align and reuse entities defined across different domains and KGs. Furthermore, reuse of (or conformance to) existing upper level ontologies, e.g., BFO (Basic Formal Ontology) [3] provides the basis for the consistent and unambiguous formal definition of entities and relations that prevents errors in coding and annotation. The alignment of ontologies in life sciences and other domains is an active area of research, and we provide an overview of recent technical developments and challenges in Section 3.3. Whilst bottom-up approaches, especially those based on alignment, are becoming more viable, especially given the support of language models, such as BERT [63], their performance is not always adequate for the task, as discussed in the second challenge in Section 3.4. Knowledge enrichment and completion improve the KG quality by performing reasoning (KG materialization), inference [56] and optimization. Reasoning and inference support the assertion of new relations based either on logical reasoning (e.g., [128, 170]) or machine learning techniques (e.g., statistical relational learning or through embedding based link predictors for new concepts [34, 35, 66, 76] and node classifiers, also called KG refinement [135]). The extent and type of logical inferences depends on the expressivity of the ontology built in phase 2, or in a bottom-up fashion in this phase, together with any associated mappings. Description Logic formalisms, such as OWL, use logic-based reasoning for detecting and correcting incorrect assertions and ontology alignments [24].
3.1.5 KG storage and inspection